OpenStudy (anonymous):

solve the equation \[2^{4x}=9^{x-1}\] and check

OpenStudy (anonymous):

start with \[4x\ln(2)=(x-1)\ln(9)\] and then do some algebra to solve for \(x\)

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

@satellite73 uhh... i feel really stupid asking this but i bring the 4x and x-1 to one side right? and also the ln9 and ln2 to one side? @Green52017 hm? lol you're welcome... ? :S

OpenStudy (anonymous):

don't forget that \(\ln(9)\) and \(\ln(2)\) are constants, treat them like \(a\) and \(b\)

OpenStudy (anonymous):

Why would you feel stupid but btw im not solving it its asking for a multiple choice im just stuff a lil bit i need a little bit of guidance lol

OpenStudy (anonymous):

A) Switch the x and y in the equation B) Change the slope of 2 to a slope of 3. C) Change the y-intercept of 2 to a y-intercept of 3. D) Change the y-intercept of 200 to a y-intercept of 3.

OpenStudy (anonymous):

so first step is to multiply out on the right, get \[4\ln(2)x=\ln(9)x-\ln(9)\] then maybe \[4\ln(2)x-\ln(9)x=-\ln(9)\] \[(4\ln(2)-\ln(9))x=-\ln(9)\] and finally \[x=\frac{-\ln(9)}{4\ln(2)-\ln(9)}\] there are other way you can write it

OpenStudy (anonymous):

how did you get 4ln(2)x ... ? ._.

OpenStudy (anonymous):

and ln(9)x

OpenStudy (anonymous):

dont forget they are constants suppose you wanted to solve \[4\times x\times A=(x-1)\times B\] for \(x\) what would you do?

OpenStudy (anonymous):

first multiply out on the right, and rewrite the left as \[4Ax=Bx-B\] then all terms with \(x\) on the left get \(4Ax-Bx=-B\) then factor out the \(x\) get \((4A-B)x=-B\) and finally divide get \[x=\frac{-B}{4A-B}\]

OpenStudy (anonymous):

replace \(A\) by \(\ln(2)\) and \(B\) by \(\ln(9)\) and you get the same thing

OpenStudy (anonymous):

okay. uhm just one last question, the sheet tells me that the asnwer is x= - log 9/log(16/9) would it be th same as ssaying x=−ln(9)/4ln(2)−ln(9)

OpenStudy (anonymous):

yes because \[4\ln(2)=\ln(2^4)=\ln(16)\]

OpenStudy (anonymous):

and \[\ln(16)-\ln(9)=\ln(\frac{16}{9})\] as i said there are other ways to write it

OpenStudy (anonymous):

YAY! okay thank you :)

OpenStudy (anonymous):

yw